# FastLGS: Speeding up Language Embedded Gaussians with Feature Grid Mapping

### v2

@@ 4.3. Ablations
This is ablation.
More visual results on different datasets. Example 1: Query relevancy visualization. Example 2&3: Object retrieval. Example 4: Downstream 3D segmentation. Example 5: Downstream 3D inpainting.
5. Conclusions and future work
In this paper, we propose FastLGS, a method that speeds up to build accurate 3D open-vocabulary semantic fields within 3DGS. By mapping and restoring multi-view CLIP embeddings through feature grids instead of MLPs, FastLGS not only has interactive query speed, but also provides more accurate target masks and semantic relevancy along with lower consumption compared with state-of-the-art methods. Meanwhile, experiments show FastLGS’s ability to be integrated with other downstream 3D manipulation tasks. While FastLGS now follows an object-centric paradigm, challenges occur in querying for parts of an object, such as “bicycle handlebars” or “chair leg”. We believe this could possibly be solved by building the feature grid in a multi-granularity way, where actual performance (time and space) along with cross-view consistency of both mask extraction and query results remain a problem, which requires further researches.

### v4

@@ Introduction
Conclusion
In this paper, we propose FastLGS, a method that speeds up to build accurate 3D open-vocabulary semantic fields within 3DGS. By mapping and restoring multi-view CLIP embeddings through feature grids instead of MLPs, FastLGS not only has interactive query speed, but also provides more accurate target masks and semantic relevancy along with lower consumption compared with state-of-the-art methods. Meanwhile, experiments show FastLGS’s ability to be integrated with other downstream 3D manipulation tasks. While FastLGS now follows an object-centric paradigm, challenges occur in querying for parts of an object, such as “bicycle handlebars” or “chair leg”. We believe this could be solved by building the feature grid in a multi-granularity way, where actual performance along with cross-view consistency of both mask extraction and query results remains a problem, which requires further research.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62302167, U23A20343, 62222602, 62176092, and 62476090); in part by Shanghai Sailing Program (23YF1410500); Natural Science Foundation of Shanghai (23ZR1420400); in part by Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission (23CGA34), in part by CCF-Tencent RAGR20240122.